For Question 1-8, consider a competitive market for a good where the demand curve is determined by the demand function: P=5-QD and the supply curve is determined by the supply function: P=QS. Where P stands for Price, QD is quantity demanded and QS is quantity supplied.

What is the equilibrium price level for the good in the competitive market?
Answer #1

Equilibrium will be attained at a price where QD=QS=Q

i.e.

5-Q=Q

2Q=5

Q=5/2=2.5 (equilibrium output)

P=5-QD=5-2.50=$2.50 (equilibrium price for good)

Equilibrium price level for the good in the competitive market will be $2.50